OpenStudy (anonymous):

A cone with a radius of 3" has a total area of 24π sq in. Find the volume of the cone.

OpenStudy (hba):

Lateral surface area of cone = π r s Surface area of base of cone = π r² where s = slant height, r = 3 SA = π r s + π r²

OpenStudy (hba):

put r =3 in above formula

OpenStudy (hba):

@alli-rose Do it ?

OpenStudy (anonymous):

um o.O i have noooo idea :/

OpenStudy (hba):

what you dont know how to plug in a certain value ?

OpenStudy (anonymous):

math is a terrible subject for me :/ im just trying to graduate and need help

OpenStudy (hba):

Well,you have to participate ..

OpenStudy (hba):



OpenStudy (anonymous):

so SA = pi 3 s + pi \[3^{2}\] or something?

OpenStudy (hba):

yeah right now see what is the s.a in your question ?

OpenStudy (anonymous):

All you have to do is find the slant height . Using the radius and slant height find the height of the cone. Then plug those values to this equation to get the volume.. \[V = \frac{1}{3}*\pi r^{2}*h\]

OpenStudy (hba):

@alli-rose

OpenStudy (hba):

SA=Surface Area

OpenStudy (anonymous):

ugh forget it -.- im so lost already thank you though

OpenStudy (hba):

No Wait @alli-rose

OpenStudy (hba):

See what is the totalarea in your question ?

OpenStudy (hba):

\[SA= 3s π + 3^2π\]

OpenStudy (anonymous):

First of all to calculate the slant height of the cone.

OpenStudy (anonymous):

Total surface area of a cone= \[3.14*radius ^{2} + 3.14*radius*slant height\]

OpenStudy (anonymous):

You have the area and the radius but you don't have the slant height so, area=75.4 radius=3 slant height=h

OpenStudy (anonymous):

75.4=\[3.14*3^{2} + 3.14*3*h\]

OpenStudy (anonymous):

to find the volume, you use : V = 4/3 π r3(cubed) so... V = 4/3 π 3(cubed) V = 4/3 π 9 V = 36 π

OpenStudy (anonymous):

so it would be 36 π in cubed .